OpenStudy (anonymous):

Classify the system as a consistent or inconsistent and dependent or independent. x+3y=17 2x+6y=34

jimthompson5910 (jim_thompson5910):

If z = x+3y, then 2z = 2x+6y So the system becomes z=17 2z = 34 Solve the second equation for z to get z = 17. This shows us that the system is consistent and dependent.

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

sure thing